# Optimizing LLM-Based Multi-Agent System with Textual Feedback: A Case Study on Software Development

### v1

@@ Abstract
[…]
Optimizing LLM-Based Multi-Agent System with Textual Feedback:
A Case Study on Software Development
Ming Shen1111thanks: Work done as an intern at AWS, Raphael Shu2222, Anurag Pratik2222, James Gung2222, Yubin Ge2222, Monica Sunkara2222, Yi Zhang2222 $1$Arizona State University, USA $2$Amazon Web Services mshen16@asu.edu {zhongzhu, anuragik, gungj, yubinge, sunkaral, yizhngn}@amazon.com
@@ 1 Introduction
Autonomous agents utilizing large language models have achieved promising results on tasks across various domains such as reasoning Yao et al. (2023); Shinn et al. (2023), code generation Shinn et al. (2023), tool usage Cai et al. (2023), and embodied AI Ahn et al. (2022). Recent studies have demonstrated that incorporating synergistic agents into multi-agent collaboration frameworks substantially further enhances NLP problem-solving abilities Zhuge et al. (2023); Liang et al. (2023); Du et al. (2024); Wu et al. (2024b); Wang et al. (2024b); Li et al. (2023a); Hao et al. (2023); Zhang et al. (2024a); Dong et al. (2024); Jiang et al. (2023); Wu et al. (2024a); Chen et al. (2024) and successfully addresses complex real-world challenges, including human behavior simulation Park et al. (2023), software development Qian et al. (2024); Hong et al. (2024), code issue resolving Tao et al. (2024), and web task execution Zhang et al. (2024b). In our study, we focus on role-based multi-agent systems, where LLM-based agents are assigned distinct roles to accomplish a task together collaboratively.
A primary constraint inherent to LLMs is their reliance on prompt design Brown et al. (2020); Reynolds and McDonell (2021); Si et al. (2023), which also extends to LLM-based agents. Specifically, it is crucial for developers to meticulously curate agent prompts, considering their respective roles and responsibilities in the collaboration framework. To this end, various automatic agent prompt optimization methods have been proposed in recent works, such as DSPy Khattab et al. (2024), GPTSwarm Zhuge et al. (2024), and TextGrad Yuksekgonul et al. (2024). However, none of these works have conclusively demonstrated the efficacy of their approaches within a genuine role-based multi-agent system, wherein agents take on various roles to tackle a complex task collaboratively. Furthermore, current works primarily focus on traditional NLP benchmarks such as MMLU Hendrycks et al. (2021a), MATH Hendrycks et al. (2021b), and HumanEval Chen et al. (2021). In contrast, our study aims to address the more intricate real-world software development task.
@@
  • We compare optimization effectiveness in various settings, including online against offline and individual against group optimization. We demonstrate online and group optimizations as more effective settings than individual and offline optimization.
@@ 2.1 LLM-Based Agents
LLM-based agents Yao et al. (2023); Shinn et al. (2023) have achieved impressive results on various NLP tasks. Applications Richards (2023) utilizing a single LLM-based agent, such as AutoGPT Richards (2023) and LangChain Chase (2022), can accomplish more functionalities and provide opportunities to a diverse audience spectrum, including developers and even non-technical users. More recently, collaborations among multiple LLM-based agents demonstrated even more capabilities. Liang et al. (2023) and Du et al. (2024) incorporate multiple LLM-based agents into a debate framework, aiming to stimulate divergent thinking and enhance both factuality and reasoning capabilities. Utilizing the unique strengths and knowledge of each individual agent, LLM-based agents can also be synergistically integrated to collaboratively enhance problem-solving capabilities on a broad range of tasks Hao et al. (2023); Zhuge et al. (2023); Zhang et al. (2024a); Dong et al. (2024); Wu et al. (2024b); Zhang et al. (2024b). In our work, we focus on role-based multi-agent system Wang et al. (2024b); Li et al. (2023a); Park et al. (2023); Qian et al. (2024); Wu et al. (2024a); Tao et al. (2024); Chen et al. (2024), in which individual agent components are assigned distinct roles to interact and collaborate to accomplish a task effectively. Among the above role-based multi-agent systems, Hong et al. (2024) and Qian et al. (2024) design agent roles inspired by Standardized Operating Procedures (SOPs) to solve software development tasks collaboratively.
@@ 2.3 Agent Optimization
We have seen various prompt optimization methods in §2.2, and it is also a critical topic for agent optimization. DSPy Khattab et al. (2024) views LLM-based systems as programs and proposes to build and optimize them in a programmatic fashion. DSPy can optimize LLM inference prompts, including few-shot examples and system instructions, through search algorithms in a combinatorial space. TextGrad Yuksekgonul et al. (2024) takes a different perspective of backpropagation being a general and powerful framework to optimize the LLM-based agent system based on natural language feedback. However, none of them proves the effectiveness of their methods in a role-based multi-agent system setting through their experimental results. Research works in another agent optimization direction, such as DyLAN Liu et al. (2024) and GPTSwarm Zhuge et al. (2024), focus on optimizing the workflow of multi-agent systems. They usually view multi-agent systems as directed graphs and optimize the graph by selecting or pruning existing edges that represent information exchange between agents. However, a fundamental limitation of such an optimization paradigm is that it relies on the assumption of the multi-agent system being decentralized, where all the agent components can draw conclusions regarding a given task, and later, the final output is generated using a consensus schema such as majority voting. This optimization paradigm is less applicable in scenarios where information flows are predefined or where agent components assume varied responsibilities to collaboratively accomplish a task, particularly in the context of real-world tasks. Finally, a significant difference between our work and all the above works is that other than traditional NLP tasks, we study optimization for software development tasks, which are complex, open-ended tasks.
@@ 3.2 Optimization Dimensions
  • Robustness – Through the robustness dimension, we aim to measure whether the software is reliable enough to handle different unexpected user inputs or other exceptions.
  • Test Case Coverage – Test cases are essential for software code. To improve software code quality and coverage, they help verify and validate whether software code is functioning as expected and identify bugs and errors in the code. We aim to optimize the software solution code to contain test cases to cover all aspects of the task description.
@@ 4.2 Optimization Pipeline
We propose a two-step pipeline to optimize a role-based multi-agent system. As shown in Figure 1, utilizing natural language feedback $I$ [Uncaptioned image] of the multi-agent system’s output [Uncaptioned image] generated by the critic mechanism $f$ [Uncaptioned image], we perform prompt optimization for participation agents inside the system. Specifically, we first use a locator $\mathcal{L}$ [Uncaptioned image] to identify the underperforming agents, [Uncaptioned image] and [Uncaptioned image], that are not performing well and generate explanations of their failures, utilizing textual feedback $I$4.2.1). Then, an optimizer $\mathcal{O}$ [Uncaptioned image] optimizes the system prompts of the identified underperforming agents utilizing explanations of their failures (§4.2.2).
@@ 4.2.1 Locating
Natural language feedback $I$ for the multi-agent system’s output provides guidance regarding which directions to improve agent components in the system. However, it is hard to improve individual agents solely based on global feedback of the final collaboration output. As a result, we design an LLM-based locator $L$ [Uncaptioned image] better capable of generating fine-grained explanations to guide optimization in the next step. Our locator takes three major components as inputs: software task description $X$, global natural language feedback $I$, and multi-agent collaboration details $G$. Collaboration details include two major components: role descriptions and communication trajectories of all participation agents. With a carefully designed prompt (see Figure 9 in Appendix), the locator focuses on the negative aspects of the global feedback and navigates to underperforming agents $N_{f}\subseteq N$ responsible for the negative aspects of feedback. The locator also generates specific failure explanations $E=\{E_{n}|n\in N_{f}\}$ of identified underperforming agents as fine-grained signals to better guide the later optimization step. Overall, the input and output workflow of the locator is $N_{f},E=L(X,I,G)$.
@@ 4.2.2 Optimizing
After identifying the underperforming agents and fine-grained explanations of their failures from the locator in the previous step, we utilize an LLM-based optimizer $O$ [Uncaptioned image] to optimize the system prompts of underperforming agents. For each underperforming agent $n$, the optimizer takes two critical components as input: input-output pair $M_{n}$ and fine-grained failure explanations $E_{n}$ of this agent. The input-output pair includes the system message $P_{n}$, all user messages, and the output of the agent. We carefully design the prompt (see Figure 10 in Appendix) to guide the optimizer $O$ to optimize the system prompt of underperforming agents in the direction where the fine-grained failure explanations can be mitigated: $P_{n}=O(P_{n},E_{n},...)$.
@@ 5.4 Baselines
0%20%40%60%80%100%66666.56.56.56.577777.57.57.57.588888.58.58.58.5Optimization Steps (%)Evaluation ScoresFunctionality ($\uparrow$)traindev0%20%40%60%80%100%6.56.56.56.577777.57.57.57.588888.58.58.58.5Optimization Steps (%)Robustness ($\uparrow$)traindev0%20%40%60%80%100%00222244446666888810101010Optimization Steps (%)Test Case Coverage ($\uparrow$)traindev0%20%40%60%80%100%0055551010101015151515202020202525252530303030Optimization Steps (%)Documentation ($\uparrow$)traindev0%20%40%60%80%100%00222244446666888810101010Optimization Steps (%)Violation ($\downarrow$)traindev
@@ 5.5 Main Results
We show the main results in Table 1. Both one-shot and direct optimization baselines cannot consistently outperform unoptimized system across all evaluation dimensions. They both fail on the functionality dimension, which is probably the most important evaluation dimension for software development. Direct optimization also fails on the code violation dimension. However, our proposed two-step optimization pipeline effectively optimizes the multi-agent system, evaluated by all optimization dimensions across all optimization settings except for only offline individual setting on the functionality dimension, where we observe only 0.1 behind the unoptimized system. Our best-performing optimization setting, online group optimization with one-pass prompting, is always better than the two baselines. Our optimization method outperforms the one-shot baseline in 22 out of the total 30 cases across all evaluation dimensions and optimization settings and always outperforms the direct optimization baseline.
@@
0%20%40%60%80%100%22224444666688881010101012121212Optimization StepsEvaluation Scores ($\downarrow$)Online Multi-PassOnline One-PassOffline Multi-PassOffline One-Pass
@@ 5.6.2 Starting from "Empty" Prompts
Instead of starting optimization from a default prompt, we study whether our optimization pipeline is still effective when the prompt to optimize starts from empty. Note that the "empty" prompts are not completely empty as they still contain very basic contexts (prompts in black) as shown in Figures 4 and 5. We slightly increase the optimization steps from the default number of 5 to 8. We analyze online group optimization with one-pass promoting across all evaluation dimensions. As shown in Table 2, starting from an "empty" prompt, our pipeline is still able to optimize the multi-agent system to a level that is on par or just slightly worse than the system optimized starting from an informative default starting prompt.
5.6.3 Case Study
We provide optimized agent prompts at the final step for functionality, robustness, and code style violation evaluation dimensions for 2 agent roles: programmer agent in Table 4 and software test engineer agent in Table 5 in Appendix. We generally observe agent prompts being optimized towards the evaluation dimension, as shown in green text. However, we also notice a current problem: as shown in the red text, the optimized prompts might contain instance-specific content that does not apply to general software development tasks even though we deliberately prompt the optimizer to think generally instead of focusing on the current task only. We leave mitigating this problem to future work.
@@ 5.6.3 Case Study
Table 2: The effect of starting optimization from an "empty" agent prompt instead of a default prompt.
@@ 6 Conclusion
Limitations
First, we conduct a case study on the group optimization of role-based LLM-based multi-agent systems on software development tasks. Thus, the scope of the study is limited to software development. A natural next step is to work on other real-world domains such as GitHub issue resolving or web task completion to which role-based multi-agent systems can be applied. Second, it could be more beneficial for us to explore other decision-making strategies in the multi-agent system. We currently use a vertical decision-making structure following previous work, as mentioned in the Appendix; however, we could explore another decision-making strategy, such as horizontal decision-making. Finally, we have only conducted experiments with OpenAI’s GPT-series model. It would also be better to try out another model family, such as Anthropic’s Claude.
@@ Appendix A Role-based Multi-Agent System
Various decision-making structures among agents Chan et al. (2024); Wu et al. (2024a) have been investigated. Following AgentVerseChen et al. (2024), we adopt the vertical decision-making structure as it is a better fit for software development. Inside the vertical structure, given the software task description $X$, a solver agent $S$ proposes a solution $Y_{t}$ at iteration $t$. Other agents, as reviewer agents $\mathcal{R}$, provide feedback =$\mathcal{F}=\{r_{t}^{i}=R_{i}(Y_{t})|R_{i}\in\mathcal{R}\}$ regarding solution $Y_{t}$ to the solver agent. Finally, the solver agent refines its solution based on the feedback $Y_{t+1}=S(X,Y_{t},\mathcal{F})$. Such a review iteration can go on for a few rounds. In the current case study, we use a total of 2 reviewer agents and a single review iteration. To determine the concrete role descriptions for solver and reviewer agents, we utilize a recruiter agent $A$ to select role descriptions for the solver and reviewer agents from a pre-defined expert pool based on the current task description. The agent role description pool we use is adapted from ChatDev Qian et al. (2024). In our study, we aim to optimize the system prompts of all agents in the vertical decision-making structure.
Appendix B Prompt for Model-Based Evaluation

### v2

@@ 1 Introduction
Autonomous agents utilizing large language models have achieved promising results on tasks across various domains such as reasoning Yao et al. (2023); Shinn et al. (2023); Ge et al. (2025), code generation Shinn et al. (2023), tool usage Cai et al. (2023), and embodied AI Ahn et al. (2022). Recent studies have demonstrated that incorporating synergistic agents into multi-agent collaboration frameworks substantially further enhances NLP problem-solving abilities Zhuge et al. (2023); Liang et al. (2023); Du et al. (2024); Wu et al. (2024b); Wang et al. (2024b); Li et al. (2023a); Hao et al. (2023); Zhang et al. (2024a); Dong et al. (2024); Jiang et al. (2023); Wu et al. (2024a); Chen et al. (2024) and successfully addresses complex real-world challenges, including human behavior simulation Park et al. (2023), software development Qian et al. (2024); Hong et al. (2024), code issue resolving Tao et al. (2024), and web task execution Zhang et al. (2024b). In our study, we focus on role-based multi-agent systems, where LLM-based agents are assigned distinct roles to accomplish a task collaboratively.
A primary constraint inherent to LLMs is their reliance on prompt design Brown et al. (2020); Reynolds & McDonell (2021); Si et al. (2023), which also extends to LLM-based agents. Specifically, it is crucial for developers to meticulously curate agent prompts, considering their respective roles and responsibilities in the collaboration framework. To this end, various automatic agent prompt optimization methods have been proposed in recent works, such as DSPy Khattab et al. (2024), GPTSwarm Zhuge et al. (2024), and TextGrad Yuksekgonul et al. (2024). However, none of these works have conclusively demonstrated the efficacy of their approaches within a genuine role-based multi-agent system, wherein agents take on various roles to tackle a complex task collaboratively. Furthermore, current works primarily focus on traditional NLP benchmarks such as MMLU Hendrycks et al. (2021a), MATH Hendrycks et al. (2021b), and HumanEval Chen et al. (2021). In contrast, our study aims to address the more intricate real-world software development task.
@@
  • We compare optimization effectiveness in various settings, including online against offline and individual against group optimization. We demonstrate online and group optimizations as more effective settings than individual and offline.
@@ 2.1 LLM-Based Agents
LLM-based agents Yao et al. (2023); Shinn et al. (2023) have achieved impressive results on various NLP tasks. Applications Richards (2023) utilizing a single LLM-based agent, such as AutoGPT Richards (2023) and LangChain Chase (2022), can accomplish more functionalities and provide opportunities to a diverse audience spectrum, including developers and even non-technical users. More recently, collaborations among multiple LLM-based agents demonstrated even more capabilities. Liang et al. (2023) and Du et al. (2024) incorporate multiple LLM-based agents into a debate framework, aiming to stimulate divergent thinking and enhance both factuality and reasoning capabilities. Utilizing the unique strengths and knowledge of each individual agent, LLM-based agents can also be synergistically integrated to collaboratively enhance problem-solving capabilities on a broad range of tasks Hao et al. (2023); Zhuge et al. (2023); Zhang et al. (2024a); Ge et al. (2023); Dong et al. (2024); Wu et al. (2024b); Zhang et al. (2024b). In our work, we focus on role-based multi-agent system Wang et al. (2024b); Li et al. (2023a); Park et al. (2023); Qian et al. (2024); Wu et al. (2024a); Tao et al. (2024); Chen et al. (2024), in which individual agent components are assigned distinct roles to interact and collaborate to accomplish a task effectively. Among the above role-based multi-agent systems, Hong et al. (2024) and Qian et al. (2024) design agent roles inspired by Standardized Operating Procedures (SOPs) to solve software development tasks collaboratively.
@@ 2.3 Agent Optimization
We have seen various prompt optimization methods in §2.2, and it is also a critical topic for agent optimization. DSPy Khattab et al. (2024) views LLM-based systems as programs and proposes to build and optimize them in a programmatic fashion. DSPy can optimize LLM inference prompts, including few-shot examples and system instructions, through search algorithms in a combinatorial space. TextGrad Yuksekgonul et al. (2024) takes a different perspective of backpropagation being a general and powerful framework to optimize the LLM-based agent system based on natural language feedback. However, none of them proves the effectiveness of their methods in a role-based multi-agent system setting through their experimental results. Research works in another agent optimization direction, such as DyLAN Liu et al. (2024) and GPTSwarm Zhuge et al. (2024), focus on optimizing the workflow of multi-agent systems. They usually view multi-agent systems as directed graphs and optimize the graph by selecting or pruning existing edges that represent information exchange between agents. However, a fundamental limitation of such an optimization paradigm is that it relies on the assumption of the multi-agent system being decentralized, where all the agent components can draw conclusions regarding a given task, and later, the final output is generated using a consensus schema such as majority voting. This optimization paradigm is less applicable in scenarios where information flows are predefined or where agent components assume varied responsibilities to accomplish a task collaboratively, particularly in the context of real-world tasks. Finally, a significant difference between our work and all the above works is that other than traditional NLP tasks, we study optimization for software development tasks, which are complex, open-ended tasks.
@@ 3.2 Optimization Dimensions
  • Robustness – Through the robustness dimension, we aim to measure whether the software is reliable enough to handle various unexpected user inputs or exceptions.
  • Test Case Coverage – Test cases are essential for software code. To improve code quality and coverage, they help verify and validate whether the code is functioning as expected and identify bugs and errors in the code. We aim to optimize the software code to contain test cases to cover all aspects of the task description.
@@ 4.2 Optimization Pipeline
We propose a two-step pipeline to optimize a role-based multi-agent system. As shown in Figure 1, utilizing natural language feedback II [Uncaptioned image] of the multi-agent system’s output [Uncaptioned image] generated by the critic mechanism $f$ [Uncaptioned image], we perform prompt optimization for participation agents in the system. Specifically, we first use a locator $\mathcal{L}$ [Uncaptioned image] to identify the underperforming agents, [Uncaptioned image] and [Uncaptioned image], that are not performing well and generate explanations of their failures, utilizing textual feedback II4.2.1). Then, an optimizer $\mathcal{O}$ [Uncaptioned image] optimizes the system prompts of the identified underperforming agents utilizing explanations of their failures (§4.2.2).
@@ 4.2.1 Locating
Natural language feedback $I$ for the multi-agent system’s output provides guidance regarding which directions to improve agents in the system. However, it is hard to improve individual agents solely based on global feedback of the final collaboration output. As a result, we design an LLM-based locator $L$ [Uncaptioned image] capable of generating fine-grained explanations to guide agent optimization. Our locator takes three major components as inputs: software task description $X$, global natural language feedback $I$, and multi-agent collaboration details $G$. Collaboration details include two major components: role descriptions and communication trajectories of all participation agents. With a carefully designed prompt (see Figure 11 in Appendix), the locator focuses on the negative aspects of the global feedback and navigates to underperforming agents $N_{f}\subseteq N$ responsible for the negative aspects of feedback. The locator also generates specific failure explanations $E=\{E_{n}|n\in N_{f}\}$ of identified underperforming agents as fine-grained signals to better guide the later optimization step. Overall, the input and output workflow of the locator is $N_{f},E=L(X,I,G)$.
@@ 4.2.2 Optimizing
[…]
@@ 5.3 Experiment Setting
5.4 Baselines
0%20%40%60%80%100%666.56.5777.57.5888.58.5Optimization Steps (%)Evaluation ScoresFunctionality ($\uparrow$)traindev0%20%40%60%80%100%6.56.5777.57.5888.58.5Optimization Steps (%)Robustness ($\uparrow$)traindev0%20%40%60%80%100%0224466881010Optimization Steps (%)Test Case Coverage ($\uparrow$)traindev0%20%40%60%80%100%05510101515202025253030Optimization Steps (%)Documentation ($\uparrow$)traindev0%20%40%60%80%100%0224466881010Optimization Steps (%)Violation ($\downarrow$)traindev
@@ 5.5 Main Results
We show the main results in Table 1. Both one-shot and direct optimization baselines cannot consistently outperform unoptimized system across all evaluation dimensions. They both fail on the functionality dimension, which is probably the most important evaluation dimension for software development. Direct optimization also fails on the code violation dimension. However, our proposed method effectively optimizes the multi-agent system evaluated on all optimization dimensions and settings except only for offline individual setting on the functionality dimension, where we observe only 0.1 behind the unoptimized system. Our best-performing optimization setting, online group optimization with one-pass prompting, is always better than the two baselines. Our method outperforms the one-shot baseline in 22 out of the total of 30 cases across all evaluation dimensions and optimization settings, and it always outperforms the direct optimization baseline.
@@ 5.6.1 Optimization Curve Analysis
0%20%40%60%80%100%2244668810101212Optimization StepsEvaluation Scores ($\downarrow$)Online Multi-PassOnline One-PassOffline Multi-PassOffline One-Pass
@@ 5.6.2 Starting from "Empty" Prompts
Figure 4: The effect of starting optimization from an "empty" agent prompt instead of a default prompt.
Instead of starting optimization from a default prompt, we study whether our optimization pipeline is still effective when the prompt to optimize starts from empty. Note that the "empty" prompts are not completely empty as they still contain very basic contexts (prompts in black) as shown in Figures 5 and 6. We slightly increase the optimization steps from the default number of 5 to 8. We analyze online group optimization with one-pass promoting across all evaluation dimensions. As shown in Table 4, starting from an "empty" prompt, our pipeline is still able to optimize the multi-agent system to a level that is on par or just slightly worse than the system optimized starting from an informative default starting prompt.
5.6.3 Case Study
We provide optimized agent prompts at the final step for functionality, robustness, and code style violation evaluation dimensions for the two agent roles: the programmer agent in Table 2 and the software test engineer agent in Table 4 in the Appendix. We generally observe agent prompts being optimized under the desired evaluation dimension, as highlighted in green text. For example, under the functionality dimension, the system prompt of the solver agent is optimized to contain "ensure that all functionalities mentioned in the task description are implemented", and system prompt of the reviewer agent is optimized to contain "provide more detailed feedback on any missing functionalities or areas for improvement". Under the dimension of code violation, the system prompt of the solver agent is optimized to contain "break down long lines of code into multiple lines to improve readability and maintainability", and the reviewer agent is optimized to contain "review the code for adherence to style guides like PEP8, including line length, and provide feedback on this aspect". In future work, we aim to improve the generalization of the optimized prompt. For example, we observe that the optimized prompt sometimes contains instance-specific content of the training instances such as "consider the lifecycle of any temporary files created during the process".
Functionality
Robustness
Violation
Table 2: We list optimized system prompts for the solver agent whose role is described as a programmer along functionality, robustness, and code violation evaluation dimensions. Green text indicates agent prompts are being optimized under the evaluation dimension.
@@ Appendix A Role-based Multi-Agent System
Various decision-making structures among agents Chan et al. (2024); Wu et al. (2024a) have been investigated. Following AgentVerseChen et al. (2024), we adopt the vertical decision-making structure as it is a better fit for software development. Inside the vertical structure, given the software task description $X$, a solver agent $S$ proposes a solution $Y_{t}$ at iteration $t$. Other agents, as reviewer agents $\mathcal{R}$, provide feedback $\mathcal{F}=\{r_{t}^{i}=R_{i}(Y_{t})|R_{i}\in\mathcal{R}\}$ regarding solution $Y_{t}$ to the solver agent. Finally, the solver agent refines its solution based on the feedback $Y_{t+1}=S(X,Y_{t},\mathcal{F})$. Such a review iteration can go on for a few rounds. In the current case study, we use a total of 2 reviewer agents and a single review iteration. To determine the concrete role descriptions for solver and reviewer agents, we utilize a recruiter agent $A$ to select role descriptions for the solver and reviewer agents from a pre-defined expert pool based on the current task description. The agent role description pool we use is adapted from ChatDev Qian et al. (2024). In our study, we aim to optimize the system prompts of all agents in the vertical decision-making structure.
@@
Dimension Name Dimension Definition
Appendix B Prompt for Model-Based Evaluation
Appendix C Are More Optimization Steps Needed?
0%10%20%30%40%50%60%70%80%90%100%666.56.5777.57.5888.58.5Evaluation ScoresFunctionality ($\uparrow$)traindev0%10%20%30%40%50%60%70%80%90%100%02244668810101212Evaluation ScoresViolation ($\downarrow$)traindev